### v2

Abstract
[…]
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
1 Introduction
Fully Homomorphic Encryption (FHE) is a cryptographic technique that enables computation directly on encrypted data without requiring decryption (10.1145/1536414.1536440). Given the extreme volume of neural network inferences processed daily (gupta2020architecturalimplicationsfacebooksdnnbased), FHE offers a solution for protecting sensitive user data in the cloud. While recent work has demonstrated FHE-based inference for CNNs and MLPs (gilad2016cryptonets; gazelle; orion), models with large embedding tables such as Deep Learning Recommendation Models (DLRMs) remain largely unexplored. As shown in Figure 1, DLRMs process sensitive user data such as demographic information, behavioral patterns, and personal preferences through embedding table lookups that transform these sparse categorical features into dense vectors (naumov2019deep).
However, a straightforward FHE approach requires encrypting massive one-hot vectors, resulting in prohibitively high computation and memory overhead. For example, the Criteo dataset contains embedding tables with over 33 million rows total, requiring thousands of FHE ciphertexts in order to perform embedding lookups for a single inference (criteo). This setup requires gigabytes of transmitted data, many expensive homomorphic operations, and latencies on the order of hours.
We develop a novel and high-performant FHE embedding table lookup optimization, which is an improvement over the prior state-of-the-art (kim2024privacypreserving). Our key insight is to develop an FHE-friendly embedding compression technique that maps original embedding tokens to deterministic coded tokens via digit decomposition that may be done client-side. We overcome several of their limitations such as underutilized ciphertext slots and a high multiplicative level consumption, while at the same time allowing for a similar exponential compression factor of embedding tables. We further propose a block-diagonal packing strategy to operate over multiple embedding tables in a parallel fashion. Our block-diagonal approach ensures that we maximize ciphertext slot utilization.
Using our FHE-friendly embedding compression technique, we present HE-LRM, a fully homomorphic solution for DLRMs to enable inference on encrypted user data without compromising privacy. We first validate our method on the UCI heart dataset and then apply our techniques to the industry-scale Criteo DLRM (heart_disease_45; criteo). While we target end-to-end encrypted DLRM inference, we note that embedding lookups are also required for transformer inference (transformer). Prior private transformer works bypass this challenge by performing embeddings client-side, exposing model parameters (nexus; thor). Our embedding solution is general and can be applied to these systems to keep embeddings server-side, improving their security.
Figure 1: Architecture of a Deep Learning Recommendation Model (DLRM). HE-LRM utilizes Fully Homomorphic Encryption (FHE) to perform end-to-end encrypted inference of this model, maintaining data privacy.
We make the following intellectual contributions:
  • HE-LRM: the first end-to-end FHE encrypted DLRM architecture supporting both dense and sparse features under FHE. We evaluate HE-LRM on two datasets, UCI Heart Disease and industry-scale Criteo.
  • An FHE-friendly embedding compression scheme that outperforms prior work kim2024privacypreserving by $77\times$ through client-side digit-decomposition. It consumes only one multiplicative level (vs. prior work’s $2+r+2s$) requiring fewer bootstraps, an expensive but necessary FHE operation for deep computations.
  • A block-diagonal embedding packing strategy to perform parallel lookups across multiple embedding tables, maximizing ciphertext slot utilization while remaining compatible with a state-of-the-art FHE linear transformation called baby-step giant-step (BSGS).
We evaluate HE-LRM on an industry-scale DLRM processing Criteo inputs, achieving end-to-end single-threaded CPU latencies ranging from 228 to 489 seconds depending upon embedding compression ratio. We also perform a case study projecting performance when FHE primitives are lowered to state-of-the-art GPU (arxiv-2024-cheddar) and custom ASIC implementations (ebel2024osiris). We find that hardware acceleration enables end-to-end latencies in the range of seconds to subseconds, making encrypted DLRM inference practical for real-world deployment.
2 Background
Figure 2: Single-threaded CPU latencies of primitive CKKS operations averaged over 30 runs. Both ciphertext-ciphertext (CT-CT) multiplication and ciphertext (CT) rotation require a compute and memory intensive key-switching operation. CKKS multiplies require ciphertexts to have at least two remaining levels.
This section provides background on the CKKS FHE scheme, the DLRM architecture, and the Orion framework for compiling PyTorch models to FHE.
2.1 Fully Homomorphic Encryption (CKKS)
We provide a systems-oriented overview of the CKKS FHE scheme. For details of the underlying cryptographic implementation, we refer the readers to the Appendix.
2.1.1 High-Level Overview
CKKS is an FHE encryption scheme that enables certain computations to be applied directly on encrypted vectors without requiring the secret key. Specifically, CKKS encrypts large floating-point vectors of a fixed power-of-two length (typically $n=2^{15}$ elements or slots) and enables SIMD addition, SIMD multiplication, and cyclic rotation upon these encrypted vectors. For addition and multiplication, one of the operands may be unencrypted (i.e., ciphertext-plaintext addition and multiplication), but the resulting output will be a ciphertext.
The atomic datatype in CKKS is a large vector of a fixed length; in other words, CKKS does not allow easy access to slicing or indexing of encrypted vectors. Despite these constraints, the CKKS instruction set (add, mul, and rotate) is sufficient to carry out complex operations such as linear layers and non-linear activation functions, making CKKS the de-facto FHE scheme for secure and outsourced deep learning applications.
2.1.2 Multiplicative Depth and Bootstrapping
Each ciphertext is associated with a multiplicative level, $\ell$, that determines the number of sequential multiplications it may undergo before decryption fails. In other words, the resulting output ciphertext of a CKKS multiplication has one less level than its inputs. When a ciphertext’s levels are exhausted, an expensive bootstrapping operation can restore levels to enable further computation. Bootstrapping is by far the most computationally expensive CKKS operation, taking roughly 20 seconds on a single-threaded CPU, and often dominates the latency of encrypted neural inference. For the cryptographic details on multiplicative depth and levels, we refer readers to Appendix A.
2.1.3 System-Level Costs
CKKS operations are both compute and memory intensive. Internally, ciphertexts are represented as integer polynomials with large coefficients (e.g., 100s to 1000s of bits) that are manipulated via modular arithmetic. As shown in Figure 2, all CKKS operations are substantially more expensive than their plaintext counterparts. Both ciphertext-ciphertext multiplication and ciphertext rotation are more expensive than all other CKKS operations due to a necessary but costly post-processing step called key-switching.
CKKS objects are also inherently large, with a single ciphertext typically around 10 to 30 MBs, and plaintext floating point vectors must be converted to plaintext CKKS objects, which also consume several MBs. Furthermore, the key-switching operation for ciphertext-ciphertext multiplication and ciphertext rotation require special evaluation keys ($\sim 200$ MB each), and a unique key is needed for each rotation offset. For this reason, most CKKS operations exceed server-grade last-level caches and require expensive DRAM accesses, further impacting performance.
2.1.4 Neural Network Layers in FHE
Linear layers such as fully connected or convolutional layers are implemented using the three CKKS primitives: SIMD addition, SIMD multiplication, and cyclic rotation. We refer the reader to Figure 2 of orion which illustrates two diagonal-based methods for performing a secure and outsourced linear transformation (i.e., a user sends their encrypted input to a server). The Halevi-Shoup method multiplies each diagonal of the plaintext weight matrix with the aligned input ciphertext, requiring $O(n)$ rotations. The Baby-Step Giant-Step (BSGS) optimization pre-rotates diagonals to reduce the rotation count to $O(\sqrt{n})$. Systems Implication: While linear layers only consume one multiplicative level, they require a large working memory for the evaluation keys for each rotation amount and for storing the plaintext weight matrix as its equivalent CKKS plaintext(s).
In contrast to linear layers, element-wise activation functions (e.g., ReLU or SiLU) must be approximated using high-degree polynomials that are comprised of SIMD multiplications and additions. The standard solution in CKKS is to use a composition of Chebyshev polynomials: for example, ReLU is commonly approximated using a composition of three polynomials of degrees 15, 15, and 27 lee2022. Systems Implication: Precise and accurate non-linear approximations necessitate high-degree polynomials which in turn have a high multiplicative depth. These approximations are the root cause of frequent bootstraps in FHE neural networks.
2.2 Deep Learning Recommendation Models
Deep Learning Recommendation Models (DLRMs) (naumov2019deep) are neural network architectures that accept as input both continuous (dense) and categorical (sparse) features and output probabilities for recommendation systems. Figure 1 shows an example of a typical DLRM: dense features are processed through an MLP whereas sparse features are transformed into dense vector representations via embedding table lookups. These processed inputs are then combined via an interaction operator after which they are passed through another MLP. The model outputs a probability corresponding to the likelihood of accepting a particular recommendation. In the context of advertising, the model predicts the likelihood of a user clicking an ad. More precisely, DLRMs consist of the following steps:
  1. 1.
    A bottom MLP processes the dense inputs. This MLP consists of a series of linear layers with ReLU activation functions.
  2. 2.
    The sparse input features are transformed into dense embedding vectors with each feature having a unique embedding table. For industry-scale DLRMs, embedding tables can contain millions of rows and consume GBs to TBs of memory (yin2021tt). These tables incur both large compute and memory costs in FHE, challenges which we address in this paper.
  3. 3.
    The dense output from the bottom MLP and the embedding vectors are combined via an interaction operation, which can either be a dot product between all pairs of embedding vectors and the dense feature or a simple concatenation of the embedding vectors and the dense feature. In this work, we opt for the latter.
  4. 4.
    The combined inputs are processed through a top MLP that outputs a single logit. This logit is passed through a final sigmoid layer to produce a probability.
While originally developed for click-through rate prediction in targeted advertising, DLRMs are flexible and can be adapted to any classification problem that contains both dense and categorical inputs. In this paper, we additionally apply the DLRM architecture to the UCI Heart Disease dataset (heart_disease_45).
2.3 The Orion Private Inference Framework
Listing 1: Top MLP in Orion.
import orion.nn as on
class TopMLP(on.Module):
def __init__(self, input_len):
super().__init__()
self.fc1 = on.Linear(input_len, 512)
self.fc2 = on.Linear(512, 256)
self.fc3 = on.Linear(256, 1)
self.r1 = on.ReLU(degrees=[15, 15, 27])
self.r2 = on.ReLU(degrees=[15, 15, 27])
self.sigmoid = on.Sigmoid(degree=31)
def forward(self, x):
x = self.r1(self.fc1(x))
x = self.r2(self.fc2(x))
x = self.sigmoid(self.fc3(x))
return x
In order to perform an end-to-end evaluation of privacy-preserving DLRM inference, we implement our compressed embedding technique and the full DLRM architecture in Orion, a PyTorch to FHE framework (orion). Orion compiles PyTorch models to FHE programs for encrypted execution, handling both linear and non-linear activations. For linear layers, Orion lowers to an optimized BSGS matrix-vector product algorithm known as double-hoisting which reduces the cost of each homomorphic rotation (dhbsgs). We find this linear transformation algorithm to be well-suited for our compressed embedding technique since we cast embedding lookups as linear transformations. For non-linear activations, Orion approximates functions using composite Chebyshev polynomials.
Furthermore, Orion automates the process of bootstrap placement and noise management within CKKS allowing us to focus on our implementation details directly in PyTorch. As an example, Listing 1 illustrates our implementation of the Top MLP in Orion. Currently, Orion lowers to a single-threaded CPU implementation of CKKS (lattigo), which we use to measure end-to-end latency in Section 6.3. However, Orion emits a static trace of each operation performed (e.g., bootstrapping, linear transformation) during inference along with the level and configuration for which they are performed. In Section 6.4, we use this static trace to estimate the latency when CKKS operations are lowered to both a GPU (arxiv-2024-cheddar) and ASIC (ebel2024osiris) backend.
3 Threat Model
We assume a semi-honest threat model  pragmaticmpc in which an adversary faithfully takes part in the private inference protocol but may try to learn additional information from the messages they receive. This threat model (also known as honest-but-curious) is also assumed in a majority of prior works in private inference gazelle; orion.
We assume a server owns the neural network and the client holds the query. We also assume that the client knows the input dimension to the neural network for which they query. In this paper, we make this assumption explicit given that the inputs are no longer only dense features but also the sparse features used for the embedding tables. To this end, we assume that the client knows the input size to the Bottom MLP and the sizes (number of rows) of each embedding table. Crucially, the client does not know the embedding dimension (number of columns) or parameters of the embedding tables, the interaction method, or the parameters of the MLPs.
4 Privacy-Preserving Embedding Lookups
4.1 One-hot solution
As explained in Section 2.2, the categorical (sparse) features are represented by embedding vectors stored in embedding tables. While a cleartext DLRM can perform standard embedding lookups directly, the homomorphic setting requires careful design of an efficient retrieval algorithm.
Embedding tables in DLRMs are parameterized as matrices $E\in\mathbb{R}^{k\times d}$ where $k$ is the number of unique categories and $d$ is the hidden or embedding dimension. For standard cleartext computations, the inputs to the embedding table are simply indices $i\in\mathbb{Z}_{k}$, which are used to retrieve rows (i.e., $E[i]$), from the embedding table for further processing by the DLRM. However, this process of extracting rows using indices is challenging in the context of CKKS where the only available operations are SIMD addition, multiplication, and cyclic rotation of encrypted vectors.
A straightforward FHE-friendly solution to this lookup problem is to treat the embedding lookup as a linear transformation by an encrypted one-hot encoded index. First, define the one-hot encoding function as $\mathsf{OHE}:\mathbb{Z}_{k}\to\{0,1\}^{k}$ such that for any $i\in\mathbb{Z}_{k}$, we have
$\mathsf{OHE}(i)=(e_{0},\dots,e_{k-1})\text{ }\text{where}\text{ }e_{j}=\begin{cases}1&\text{if }j=i\\ 0&\text{otherwise}\end{cases}$
Figure 3: Embedding table sizes (number of rows) for each of the 26 categorical features in the Criteo dataset for a total of $33.8$ million rows.
Figure 4: Prior work kim2024privacypreserving compressed embedding lookup with $k$ rows and an embedding dimension of size $d=3$. The client must store the coded token mapping of size $k\times\ell$ locally. Prior work performs this compressed lookup homomorphically by utilizing $p\ell$ slots per ciphertext and performing one homomorphic $\mathsf{Indicator}$ function followed by $d$ multiplications with the concatenated columns of the compressed tables followed by $d\log(p\ell)$ rotations and summations (see Table 4 of kim2024privacypreserving). Rotation and summation within a ciphertext produces several wasted slots that are filled with invalid data. In this example, two separate tokens (“cat” and “dog”) fit into a single 16-slotted ciphertext.
Then, by encrypting this basis vector $\mathsf{OHE}(i)$ under FHE, the server can perform the vector-matrix multiplication $\mathsf{OHE}(i)\cdot E$ homomorphically. This encrypted computation would be equivalent to the cleartext index lookup: $E[i]$.
The primary drawback of this method is the large amount of slots required: each one-hot encoding requires $k$ slots of a CKKS ciphertext where only a single element within these $k$ slots is non-zero. Figure 3 shows that realistic DLRMs exhibit large embedding table sizes. For a DLRM tailored to the Criteo dataset, the naïve, one-hot encoding for each embedding table would require $33.8$ million slots. And for standard FHE parameters with 128-bit security with a slot count of $2^{15}$, the sparse input for Criteo would require approximately $1000$ CKKS ciphertexts.
4.2 Prior work: CodedHELUT
To avoid requiring $k$ slots for an embedding table containing $k$ rows, the prior work kim2024privacypreserving uses compressed embedding tables learning through Deep Compositional Code Learning  shu2018compressing. In this setup, a $k$-sized embedding table is decomposed into $\ell$-many, $p$-sized tables such that $k=p^{\ell}$. In this way, an index $i\in\mathbb{Z}_{k}$ is first mapped to a sequence of $\ell$ tokens $(i^{0},\ldots,i^{\ell-1})\in\mathbb{Z}_{p}^{\ell}$ and each coded token retrieves a row from each of the smaller tables. Finally, the output embeddings from each of the $\ell$ tables is summed to produce the final output embedding vector. This compression technique reduces a table by an exponential factor (from $k=p^{\ell}$ rows to only $p\ell$ rows) and experiments over the GlovE and GPT-2/Bert embedding tables display its efficacy in reducing table sizes (see Section 5 in kim2024privacypreserving).
Figure 5: Comparison of encrypted embedding lookups using a fixed set of FHE parameters for different settings of $k=p^{\ell}$ and embedding dimensions, $d$. CodedHELUT first uses the Encrypted Indicator Function to perform the one-hot encoding server side. This function requires a bootstrap operation given our FHE configuration. For the embedding lookup stage, swapping out their suggested TableMult algorithm with a double-hoisted BSGS linear transformation reduces runtime as TableMult requires separate rotations for every hidden dimension, $d$. Our proposed solution performs the one-hot encoding client-side while still requiring the same number of slots as the prior method and directly leverages double-hoisted BSGS.
The prior work of  kim2024privacypreserving leverages this compression technique, but performs the one-hot encoding server-side. They do this by building an encrypted indicator function (EIF), which effectively transforms a sequence of coded tokens into their one-hot equivalent. The EIF is a composition of two steps: a squaring method that approximates a standard indicator function $\delta_{a}:\mathbb{Z}_{p}\rightarrow\{0,1\}$ for $a\in\mathbb{Z}_{p}$, and a cleaning function applied to the output of the prior step to efficiently round the value to 0 or 1.
Given the aforementioned EIF,  kim2024privacypreserving provides two algorithms (with and without compressed embeddings) for LUT evaluation with encrypted data. The full algorithm from  kim2024privacypreserving is shown in Figure 4. We refer to Algorithm 3 and Algorithm 4 from kim2024privacypreserving for more details.
4.3 Limitations of CodedHELUT
We now detail the three primary limitations of the CodedHELUT method presented in  kim2024privacypreserving. We address each limitation in Section 5.
Slot Utilization: The CodedHELUT method requires $O(\log{p\ell})$ rotations for each output ciphertext to compute the desired embedding vector as shown in Figure 4. Their approach uses the standard logarithm-based rotation-and-summation, which we call the Rot-Sum algorithm (see Figure 4 of HECO  heco for a diagram of Rot-Sum). There are two issues with using the Rot-Sum method. First, the slot occupancy for $p\ell$ must necessarily be inflated to the closest power of 2. As an example, summing 6 slots with Rot-Sum requires padding up to 8 slots with two zeros and applying Rot-Sum to all $8$ slots. Second, a consequence of applying Rot-Sum within a ciphertext is that most slots are filled with undesired partial sums or invalid data. We visualize this in Figure 4 where, for every $p\ell=8$ slots, there are $p\ell-1=7$ slots that contain invalid data. In contrast, our BSGS-based embedding lookup supports arbitrary dimensions and does not waste any slots by summing the coded tables across ciphertexts rather than within ciphertexts.
Output ordering: While the CodedHELUT algorithm produces the correct embedding, this particular embedding vector is sharded across $d$ ciphertexts where $d$ is the embedding vector dimension. This can be seen in Figure 4 where the desired embedding vector (e.g., the vector $\left[29,31,33\right]$) is fragmented across the three output ciphertexts. In order to perform the subsequent concatenation or linear transformation in either DLRMs or LLMs, this fragmented embedding vector must be consolidated into contiguous slots. Performing this consolidation requires 1) further rotations to align the embedding elements and 2) masking to remove the junk data from the wasted slots, which increases level consumption. On the other hand, our BSGS-based method produces the correct embedding vectors in contiguous slots while consuming only a single level.
Level Consumption: Beyond the levels consumed to consolidate the output embedding vector, the indicator function incurs a multiplicative depth of $2+r+2s$ (see Section 3.3 of kim2024privacypreserving). Based on both the FHE parameters and the iteration parameters $r$ and $s$, the indicator function may require an expensive bootstrap operation. Indeed for a 128-bit secure FHE configuration, we find that we must judiciously set both the ciphertext and auxiliary moduli to ensure that the Indicator function does not overflow the available levels. Conversely, our embedded lookup only consumes one level regardless of the embedding size.
5 Our solution
We now describe our embedding compression technique, which we use to build an end-to-end FHE-friendly DLRM architecture, referred to as HE-LRM. Each HE-LRM inference takes as input an encryption of the concatenated dense and sparse features and returns an encrypted, un-normalized logit to the client.
5.1 Digit Decomposition for Compressed Tables
Recall that prior work kim2024privacypreserving built their CodedHELUT algorithm to integrate with compositional embedding tables that map from input tokens $i\in\mathbb{Z}_{k}$ to a sequence of tokens in $\mathbb{Z}_{p}^{\ell}$, and this mapping is stored as its own lookup table on the client device. While we opt for the same mapping space (i.e., $\mathbb{Z}_{k}\rightarrow\mathbb{Z}^{\ell}_{p}$), we choose a much simpler compression technique that can be seen as a generalization of the Quotient-Remainder (QR) method from shi2020compositional. The QR method decomposes a single $k$-sized embedding table into two tables: a Quotient and Remainder table by computing $(i\bmod q,i\mathbin{\%}q)$ for some $q<k$. The parameters of the QR tables are then trained end-to-end.
We extend the QR method by performing a digit decomposition of an original input token $i\in\mathbb{Z}_{k}$ into base $p$ to form a tuple of tokens in $\mathbb{Z}_{p}^{\ell}$. This process produces a maximum of $\ell=\lceil\log_{p}(k)\rceil$ tokens and allows a similar exponential compression factor of an embedding table of size $k\times d$ into $\ell$ embedding tables, each of size $p\times d$. Figure 6 (left) depicts this digit decomposition technique.
Figure 6: Our embedding lookup technique. Left: Client-side base-$p$ digit decomposition maps token “cat” (14) to coded tokens $(x_{0}=2,x_{1}=3)$. Right: Block-diagonal packing of embedding tables $E_{0}$ (compressed) and $E_{1}$ (uncompressed) enables simultaneous lookup using a single homomorphic matrix-vector product.
Crucially, our compression technique does not require learning or storing the mapping from an index to a set of coded tokens. Rather, our mapping is a deterministic function that can be computed entirely by the client given that they know the size of the compressed embedding tables, which is encompassed by our threat model. Performing this one-hot encoding prior to encryption allows us to forgo the encrypted indicator function and directly treat the embedding layer as a linear transformation which only consumes one multiplicative level. In Figure 5, we compare the CodedHELUT algorithm with our solution in terms of latency for various hidden dimensions, showing that our method outperforms CodedHELUT for all considered hidden dimensions with up to a $77\times$ speedup for the largest dimension of 768.
We train DLRM models on the Criteo Kaggle Dataset (https://huggingface.co/datasets/criteo/CriteoClickLogs) which, as seen by Figure 3, has a maximum embedding table size of approximately 10 million rows. For each training run, we set a maximum embedding table size; any embedding table larger than this threshold are compressed using our digit-decomposition technique by selecting a base $p$. We present our loss curves, test AUC measurements, and FHE latencies in Section 6.
5.2 Parallel Embeddings
Our discussion in the previous sections was limited to performing an encrypted lookup into a single (compressed or uncompressed) embedding table. However, state-of-the-art DLRMs usually contain more than one sparse feature given the problem specification. Indeed, the UCI Healthcare dataset heart_disease_45 and the Criteo Kaggle dataset which we target require 8 and 26 embedding tables, respectively. Furthermore, each embedding table has a different number of rows, and this is the case for both datasets. Under the constraints of FHE, we are posed with the following question: how can we efficiently perform encrypted lookups when we have multiple embedding tables?
A straightforward solution to this multi-embedding table case is to simply encrypt each compressed index separately and perform the methods discussed before. However, this method would require one CKKS ciphertext per embedding table and may be wasteful when compressed tables have a total size less than the number of available CKKS slots. For our use cases with both UCI and Criteo, we find that an aggressive compression ratio induces this exact scenario.
To this end, we propose a multi-embedding packing strategy which places unique embedding tables diagonally across a standard weight matrix by starting each subsequent embedding table at the bottom right corner of the previous embedding table. This packing strategy is akin to calling: torch.block_diag(*list_of_embeddings) (and indeed, we use this exact command in our implementation). An example of our packing strategy can be seen in Figure 6 where exactly two unique feature embedding tables are stored in a single weight matrix. The client simply one-hot encodes each index for each table (again, either compressed or uncompressed) and concatenates these one-hot vectors into a single vector, which can then be encrypted. The resulting matrix-vector product exactly extracts and combines the correct columns of the packed embedding matrix and places the encrypted embedding vectors into contiguous slots within an output ciphertext.
Importantly, this diagonal multi-embedding packing strategy aligns well with diagonal-based BSGS matrix-vector product strategies discussed in Section 2.3. Moreover, diagonally aligning embedding tables incurs little FHE overhead as long as the sum of the embedding dimensions does not exceed the slot count.
6 Results
In this section, we report our results of running end-to-end encrypted DLRMs directly in the Orion FHE framework. We refer to Appendix B for an illustration of the full HE-LRM architecture. We discuss our finding for training FHE-amenable DLRMs and running end-to-end FHE inference.
6.1 Setup
We conduct all of our experiments on an Intel Xeon Gold 5218 processor running at 2.30GHz with 64 CPU cores and 512 GB of RAM. We find this amount of RAM necessary to run our largest compressed DLRM models under FHE. We train all DLRMs on NVIDIA 3090 GPUs, and experimental results (training losses, AUCs, and FHE latencies) are all averaged over 3 runs. We use an A100 for Cheddar experiments (arxiv-2024-cheddar).
Figure 7: Training curves and performance trade-offs for compressed DLRM models across different compression ratios.
Our implementation builds upon the CAFE framework (https://github.com/HugoZHL/CAFE), which provides an optimized DLRM implementation. We implement the generalized QR decomposition technique within CAFE. We use the hyperparameters set by their repository. We port all DLRM models directly into the Orion framework and implement all necessary components (extraction, concatenation, embeddings). Unless stated otherwise, we choose an FHE parameter set that enables bootstrapping while maintaining 128-bit security. We use the Cheddar codebase to collect GPU performance.
Figure 8: End-to-end (aggregate) FHE latency of a single sample passed through our compressed DLRM models averaged over three runs. Exact compression ratios are listed in Table 1 in Appendix C.
We evaluate HE-LRMs on two datasets: the UCI Heart Disease (id=45) and Criteo 7-day click-through datasets. The UCI Heart Disease dataset is used to predict the detection of heart disease based on 13 attributes (5 dense and 8 sparse). Features include age, resting blood pressure, and maximum achievable heart rate. This dataset serves as a basis for testing the validity of our methodology as well as latency of small DLRMs, in a privacy-sensitive setting. The Criteo dataset includes 13 continuous features (normalized integers) and 26 categorical features where each feature has been anonymized. There are approximately 45 million training samples and this dataset is used to predict the click-through-rate for a provided advertisement. The first 6 days of data are used for training while the seventh day serves as the validation set. We are able to train a DLRM on the Criteo dataset from scratch in about 2 hours on a single 3090 GPU.
6.2 Training DLRMs
To enable efficient FHE operations on large embeddings, we implement our digit decomposition strategy (with base $p=4$) in the CAFE framework for Criteo. We set an embedding table size threshold: tables exceeding this threshold are compressed, while smaller ones remain full.
Following prior work, we define Compression Ratio (CR) as the ratio of total rows in the uncompressed DLRM to rows in the compressed model. We vary thresholds across orders of magnitude; an aggressive 500-row threshold compresses most of Criteo’s 26 sparse features, achieving $31180\times$ compression.
Training results show expected behavior: more compressed models exhibit slightly higher loss throughout training, with all configurations following similar dynamics and diurnal patterns (from 6-day in-order processing). We also train models with SiLU activation instead of ReLU, as SiLU’s smoothness enables lower-degree polynomial approximation in FHE. However, SiLU models consistently show lower test AUC and higher training loss across all configurations, revealing a latency-performance tradeoff: SiLU reduces FHE latency but sacrifices out-of-distribution performance.
6.3 End-to-End Results with Orion
UCI Heart Disease: We first apply our HE-LRM architecture to the underlying DLRM model trained for the UCI Heart Disease dataset. This dataset has 5 dense features and 8 sparse features with sizes $[2,4,2,3,2,3,4,3]$ giving an input ciphertext where 28 slots are needed. Given the size of the input, we scale down the DLRM model and use the $x^{2}$ activation function rather than ReLU. Nonetheless, this model achieves a validation accuracy of 85% and serves as a useful testing framework for our implementation.
We are able to fit all embedding tables into a single matrix-vector product by using our multi-embedding diagonal packing technique described in Section 5.2. Averaging over three runs, we achieve an end-to-end latency of 24.22 seconds for this model. Given our FHE parameter set, this model requires exactly one bootstrap taking 75% of the runtime.
Criteo: Figure 8 presents our key end-to-end FHE latency results for a single Criteo input across our compressed DLRM threshold configurations averaged over three runs. We highlight two key observations from our results.
Observation 1: All highly compressed models exhibit similar FHE latencies. The three most aggressively compressed models ($\approx 10^{3}\times$ to $10^{5}\times$ compression ratio) all exhibit similar latencies around 230 seconds. This similarity is due to two factors. First, our multi-embedding parallel packing scheme allows us to perform a single-ciphertext linear transformation for two smallest models. Both $10^{4}\times$ and $10^{5}\times$ have 512 non-zero diagonals in the plaintext blocked embedding matrix whereas the $10^{3}\times$ model has 1024 non-zero diagonals and requires two ciphertexts to hold the one-hot inputs. This means our embedding lookup takes only 3-5 seconds for these three highly compressed models. Second, apart from the embedding size, the DLRM architecture is the same for all models, meaning that all FHE operations are identical outside of the embedding lookup. Indeed, Orion places the bootstrap operations in the same locations for all three networks.
On the other hand, the larger models ($10^{2}\times$ and $10\times$) require many ciphertexts to store all one-hot coded tokens. Table 1 shows the total number of slots required by these models: $10^{2}\times$ and $10\times$ require 18 and 85 ciphertexts to hold all one-hot encoded inputs respectively. This inflates the matrix-vector product latency; for example the largest model takes 213.9 seconds to perform the embedding lookup.
Observation 2: Changing the activation function impacts the FHE level management policy. In more detail, the SiLU activation is a smoother function when compared to ReLU and can be approximated by a lower degree polynomial. This lower degree approximation, in turn, requires overall less bootstraps for HE-LRM and affects the overall level management policy output by Orion. Concretely, the SiLU based model requires just 5 bootstrap operations whereas the ReLU based models require 12 bootstraps. Furthermore, the level management policy constructed by Orion performs the embedding layers in the SiLU networks at $\ell=3$ whereas the embedding is performed at $\ell=4$ for the ReLU models. For this reason, embedding layers in the $10^{2}\times$ model take only 44 seconds for the SiLU-based DLRM, but take 51 seconds for the ReLU-based DLRM.
6.4 Case Study: Impact of Hardware Acceleration
Figure 9: Projected GPU (arxiv-2024-cheddar) and ASIC (ebel2024osiris) for the SiLU-based DLRMs.
We saw that end-to-end industry scale encrypted DLRMs require approximately 8 minutes for a single-threaded CPU FHE implementation, far from practical. In this section, we conduct a case study of lowering FHE operations to both GPU and ASIC backends. Using Orion’s static trace to run micro-kernels from Cheddar’s codebase (arxiv-2024-cheddar) and Osiris’s cost model (ebel2024osiris) for latency estimation, we find that even GPU acceleration provides roughly $200\times$ speedup over CPU, bringing previously impractical runtimes from minutes down to seconds. We examine these speedups across different compression ratios to understand where each hardware backend excels as shown in Figure 9.
In particular, the least compressed model ($10\times$) takes roughly 1.07 seconds on the GPU backend, while the ASIC maintains sub-second latency at approximately 0.037 seconds. This performance gap at larger model scales indicates where each backend excels: DLRMs that are robust to aggressive compression can achieve practical latencies on GPUs, while models requiring minimal compression to preserve accuracy necessitate ASIC acceleration for production-scale deployment.
7 Related Works
FHE Compilers.
Compiler toolchains for FHE can be categorized into two categories: circuit-level compilers heco; lee2022hecate; cowan2021porcupine; malik2023coyote; lee2023elasm, focusing on low-level optimization and scheduling for general-purpose programs, and domain-specific compilers (early works include dathathri2019chet; dathathri2020eva). Subsequent frameworks such as nGraph-HE boemer2019ngraph; boemer2019ngraph2, TenSEAL benaissa2021tenseal, SEALion van2019sealion and Concrete-ML ConcreteML provide python-based interfaces to run encrypted inference for simple classifiers or quantized networks. More recent compilers such as Dacapo cheon2024dacapo, HeLayers aharoni2020helayers, Fhelipe krastev2024tensor and Orion orion improve upon these works by introducing automating bootstrap placement and support for deeper models.
Private Inference.
In a client–server model, the model weights remain hidden from clients and inputs from the server. Private inference protocols commonly use homomorphic encryption and evaluate the whole network server-side, approximating nonlinear layers with low-degree polynomials, trading some accuracy for performance. Early systems (CryptoNets, MiniONN, GAZELLE, DiNN) demonstrated feasibility gilad2016cryptonets; liu2017oblivious; juvekar2018gazelle; bourse2018fast. For more complex tasks such as natural language processing or recommendation systems, embedding layers with large LUTs are a central challenge. In CKKS, prior work avoids large encrypted LUTs by (i) sending high-dimensional encrypted one-hots al2020privft or (ii) computing embeddings in plaintext on the client lee2022privacy; chen2022x. Both solutions have limitations with proprietary tables or tight bandwidth. Kim et al. kim2024privacypreserving enable efficient encrypted evaluation of large LUTs in CKKS with compression (see section 7.2).
7.1 FHE Accelerators
GPUs: One of the first to port FHE to GPU was cuHE cuHE, paving the way for future work such as over 100x (gpuboot) and Cheddar (arxiv-2024-cheddar).
FPGAs: Prior work heax; agrawal2022fab; poseidon employ FPGAs as lower power, more custom alternative to GPUs, and further accelerates FHE. Notably, FAB agrawal2022fab is the first to support CKKS bootstrapping on FPGAs and also employs the double-hoisting optimizations bossuat.
Custon ASICs: F1 feldmann2021f1 and Cheetah reagen2020cheetah were among the first ASIC designs for homomorphic encryption. Since then, works such as CraterLake craterlake, BTS bts, ARK ark, SHARP sharp, VPU 10992987, and Osiris (ebel2024osiris) have targeted fully homomorphic encryption, explicitly accelerating bootstrapping.
7.2 Compression techniques for DLRMs
Embedding tables in DLRMs can dominate memory and bandwidth, creating bottlenecks for training and inference. Prior work compresses these tables to cut storage and compute while preserving accuracy.
Quantization-based. Lower-precision data types reduce memory footprint through methods like uniform or row-wise quantization, applied during training (quantization-aware training, QAT) zhou2024dqrm; kang2020learning; ko2021mascot or post training guan2019post; deng2021low. Clustering-based methods (known as codebook methods) can also represent embeddings with shared centroids for high compression shu2018compressing.
Hashing-based. The Hashing Trick maps large input IDs into a smaller embedding space, trading memory for collision risk weinberger2009feature. The Quotient–Remainder trick introduced by Facebook mitigates collisions shi2020compositional. Hashing is often combined with other aforementioned methods tsang2023clustering; ghaemmaghami2022learning.
Adaptive methods. Adaptive compression methods dynamically allocate memory based on feature importance. One example is CAFE+ liu2025cafe+ which tracks importance of each feature and combines this dynamic strategy with hashing and quantization to scale to very large models.
Decomposition. Factorizations (Tucker, TT, CP) express tables in structured low-rank forms. For DLRMs, tensor-train decomposes large tables into sequences of small matrices yin2021tt.
8 Conclusion
In this paper, we presented HE-LRM, the first end-to-end implementation of a private Deep Learning Recommendation Model using Fully Homomorphic Encryption (FHE). To address the challenge of efficiently processing sparse categorical features under FHE, we developed an FHE-friendly embedding compression technique that achieves up to a $77\times$ speedup over prior work, along with a multi-embedding packing strategy that enables parallel lookup operations. We demonstrated HE-LRM on the UCI Heart Disease and Criteo Kaggle datasets, achieving private inference in 24 seconds and 227 seconds respectively on a single-threaded CPU implementation. Through performance projections on GPU and ASIC FHE backends, we show that HE-LRM latencies can be reduced to seconds and sub-seconds respectively, demonstrating that privacy-preserving recommendation systems are approaching practical deployment.
Acknowledgements
This work was supported in part by Graduate Assistance in Areas of National Need (GAANN). The research was developed with funding from the NSF CAREER award #2340137 and DARPA, under the Data Protection in Virtual Environments (DPRIVE) program, contract HR0011-21-9-0003. Reagen and Ebel received a gift award from Google. The views, opinions, and/or findings expressed are those of the authors and do not necessarily reflect the views of sponsors.
Appendix A CKKS operations
CKKS cheon2017homomorphic; ckks2 is a SIMD style homomorphic encryption scheme that encrypts a vector of complex or real values into a ciphertext. This scheme relies on RingLWE ciphertexts in $\mathcal{R}_{Q}^{2}$} for a given ring $\mathcal{R}=\mathbb{Z}[X]/(X^{N}+1)$, where $N=2^{k}$, for $k>0$, is the ring dimension and $\mathcal{R}_{Q}=\mathcal{R}/Q\mathcal{R}$ describes the ring $\mathcal{R}$ reduced modulo an integer $Q=\prod_{i=0}^{L}q_{i}$. The integer $L$ is known as the multiplicative depth and represents the maximum number of rescaling levels available before decryption fails. CKKS supports element-wise operations such as addition and multiplication as well as cyclic rotation. Based on the characteristics of the CKKS scheme, it becomes a natural choice for applications such as deep learning recommendation models where real-valued vectors are considered. We now describe some of the core operations used with CKKS.
Encoding. Consider a real (or complex) vector $\textbf{x}\in\mathbb{C}^{N/2}$. This vector can be encoded into elements of $\mathcal{R}_{Q}$ using an approximate inverse of a scaled complex canonical embedding. More precisely, one applies an inverse Fast Fourier Transform on the elements of x and scales each output by a scaling factor $\Delta$. Finally, each element is rounded to the nearest integer as encryption is performed over integers modulo $Q$. This step outputs a plaintext polynomial $\textbf{m}(X)$ which packs $N/2$ complex values. These $N/2$ values are referred to as the available slots in a CKKS object.
Encryption The plaintext polynomial $\textbf{m}(X)$ can be encrypted into a ciphertext $(\textbf{a},\textbf{b})\in\mathcal{R}^{2}_{Q}$ with a given public key and the addition of some random noise.
Addition. CKKS supports element-wise plaintext-ciphertext and ciphertext-ciphertext additions. The resulting ciphertext after this operation corresponds to the SIMD addition of the underlying complex vectors.
Multiplication. Similar to addition, CKKS supports multiplication between either a plaintext and ciphertext or between two ciphertexts. Multiplication is followed by a rescaling procedure to avoid an exponential growth in the scaling factor. The resulting polynomial after this operation has coefficients in $\mathbb{Z}_{Q_{\ell-1}}$ instead of $\mathbb{Z}_{Q_{\ell}}$ where $Q_{\ell}=\prod_{i=0}^{\ell}q_{i}$ for $0<\ell\leq L$ which corresponds to the consumption of a level in the chain of moduli.
Rotation. Cyclic rotations shift the elements of the input vector x by $0<k<N/2$ slots. The resulting ciphertext after this operation corresponds to the same underlying vector with shifted elements by $k$ slots.
Key-switching. Key-switching is a standard operation in CKKS which converts a ciphertext encrypted under one secret key into a ciphertext that decrypts correctly under another secret key. This procedure is necessary after certain operations such as ciphertext-ciphertext multiplication or ciphertext rotation so that the result remains compatible with the decryption key. The transformation is done using a special switching key, which allows this change without decrypting the ciphertext.
Bootstrapping. In CKKS, a ciphertext can be bootstrapped to increase its number of available levels. This operation is necessary after the moduli chain has been depleted ($\ell=0)$ via multiplications and no more levels are available for further computations. Among all CKKS operations, bootstrapping is the most computationally expensive, taking roughly 20 seconds on a single-threaded CPU.
Appendix B The HE-LRM architecture
Figure 10 illustrates the entire pipeline of our HE-LRM architecture.
Figure 10: The HE-LRM architecture. Beginning from the left, the client first records both their dense and sparse features for some outsourced recommendation system (e.g., a medical inquiry). The client then one-hot encodes (either compressed or uncompressed) their sparse features, appends these one-hot vectors to the dense features, and encrypts the resulting vector. The server performs the appropriate homomorphic operations to extract the dense and sparse features into two separate ciphertexts. The dense features are processed through the Bottom MLP whereas the one-hot sparse features are used as input to the diagonally-packed embedding tables in order to perform parallel embedding table lookups. The resulting vectors are concatenated and are further processed through a Top MLP before sending the encrypted logit back to the client. We implement this entire process in the Orion framework.
Appendix C Compression ratios
Table 1 reports the compression ratios, slots required and resource requirements for different embedding table thresholds.
Table 1: Compression ratio and resource requirements for different embedding table thresholds.
[…]
$500$ $31180\times$ $1096$ $64$
$5000$ $3746\times$ $9027$ $70$
$50000$ $707.1\times$ $47759$ $74$
$500000$ $59.28\times$ $569545$ $140$
$5000000$ $12.18\times$ $2772109$ $320$